OpenStudy (anonymous):

Part 1: Decide whether you would use the graphing, substitution, or elimination method to solve the following system of equations. Explain, in complete sentences, why you chose that method. Part 2: Solve the following system of equations and show all of your work. 4x + 3y = -1 3x + y = 3

OpenStudy (anonymous):

Elimination. Because, multiplying second equation by -3 allows you to eliminate y and find x. Solution: 4x + 3y = -1 3x + y = 3 4x + 3y = -1 -9x - 3y = -9 Adding: -5x = -10 => x = -10/-5 = 2 Substituting for x = 2 in equation #2, you get 3*2 + y = 3 => y = 3-6 => y = -3